please help with this word problem

How many gallons of a 5% salt solution must be mixed with 60 gallons of a 9% solution to obtain a 7 % solution?

paint costing $ 18 per gallon is to be mixed with $ 4-per gallon thinner to make 14 gallons of a paint that can be sold for $12 per gallon. How much paint thinner should be used?

Answer by checkley77(12844)

.05x+.09*60=.07(x+60)
.05x+5.4=.07x+4.2
.05x-.07x=4.2-5.4
.02x=-1.2
x=-1.2/-.02
x=60 gallons of 5& salt is needed.
Proof:
.05*60+.09*60=.07(60+60)
3+5.4=.07*120
8.4=8.4